# ValueScope: Unveiling Implicit Norms and Values via Return Potential Model of Social Interactions

### v1

@@ Appendix D Grounding 5-point Scale for Normness Ratings
In §4.2, we employ a 5-point Likert scale using GPT-3.5 to rate comments and sample them to gauge their normness. Additionally, in §4.2.2, we generate five variations of each original seed comment based on the 5 different scales of normness. Thus, for each norm dimension, we created a 5-point Likert scale and grounded their definitions in prior works Dementieva et al. (2022); Wulczyn et al. (2017); Joshi et al. (2017); Goffman (1955); Brown and Levinson (1987); Lakoff (1973). For example, we define formality based on using abbreviations, slang, colloquialisms, non-standard capitalizations, complete sentences, contractions, punctuations, and opening expressions of sentences Dementieva et al. (2022). Meanwhile, we define politeness as a set of strategies for conducting face-threatening acts while minimizing the chance that we or others will lose our positive or negative faces. Brown and Levinson (1987). The 5-point Likert scale across the norm dimensions can be found in Figures 37-39 as well as Figure 10.
Appendix E Human Annotation
@@ E.1 Normness Scale Annotation
To evaluate the NSP models and the quality of GPT-4 generated labels for student models, we curate a high-quality human annotation set of 450 samples for each norm dimension. Each sample was annotated by three volunteer annotators, who are graduate students in NLP and Linguistics at a US-based institution. We did not provide payment, but we obtained their consent to use their annotations for AI model evaluation.
@@
Across the four topics, we collected human annotations for 450 samples [^10] For all topics except “Gender,” we annotated 100 randomly-sampled pairwise comments. For “Gender” topic, we annotated 150 pairwise comments, in which 100 pairwise comments came from r/askmen and r/askwomen while the remaining 50 pairwise comments came from comparisons with one of the gender subreddits (including r/asktransgender) and r/asktransgender.. Each sample was annotated for five norm dimensions, resulting in a total of 2,250 annotations per human annotator. The average inter-annotator agreement, measured by Fleiss’s kappa, was 0.56, considered a moderate to substantial agreement Landis and Koch (1977). See Table 6 for the full agreement scores across 4 topics and 5 norm dimensions.
@@
Topic Completely Share Roughly Mostly Completely
Dissimilar Details Equiv. Equiv. Equiv.
Gender 0.02 0.16 0.16 0.51 0.16
Politics 0.02 0.16 0.29 0.4 0.13
Science 0.04 0.07 0.13 0.42 0.33
Finance 0.03 0.1 0.12 0.38 0.37
Total 0.03 0.12 0.17 0.43 0.26
Topic Not at all Somewhat Very
Gender 0.04 0.16 0.8
Politics 0.04 0.09 0.87
Science 0 0.16 0.84
Finance 0.07 0.17 0.77
Total 0.04 0.14 0.82
Topic High Quality Not High Quality
Gender 0.67 0.33
Politics 0.58 0.42
Science 0.91 0.09
Finance 0.7 0.3
Total 0.71 0.29
Topic Original Comments Synthetic Comments
Gender 0.96 0.43
Politics 0.73 0.13
Science 0.75 0.78
Finance 0.42 0.78
Total 0.81 0.5
E.2 Community Language Simulation Filter Annotation
Appendix F GPT Evaluations
Recall in §4.2.1 that we employ GPT-3.5 to sample and rate comments on a 5-point Likert scale (defined in Appendix D) for a particular norm dimension and subsequently use GPT-4 to generate binary synthetic labels comparing a pair of comments. Here, we describe our evaluation process for GPT-3.5 and GPT-4 using the human annotations from Appendix E.1.
F.1 Evaluating the Quality of GPT-3.5 Rating
@@ Appendix F GPT Evaluations
F.2 GPT-4 Automatic Pairwise Labeling
We underwent extensive prompt-tuning efforts to generate high-quality and accurate binary synthetic labels using GPT-4. Below, we discuss our prompt design choices (§F.2.1), the prompt tuning results to select the best prompt for our task (§F.2.2), and the full evaluation results of the chosen prompt against human annotations (§F.3).
F.2.1 Prompt Design Considerations
@@
F.2.2 Prompt-Tuning Results
Based on the proposed features in §F.2.1, we design multiple prompting pipelines and evaluate their performance on the binary labeling task—given two comments, compare the comments in each of the five norm dimensions. Performance is measured by the label accuracy against a human-annotated gold data, thus assessing the effect of different prompting pipelines to produce accurate labels.
@@ F.2.2 Prompt-Tuning Results
F.3 Evaluating the Chosen GPT-4 Labeling Pipeline
@@ G.3 NSP Evaluation Results
Appendix H Community Preference Prediction
H.1 CPP Training Details
@@
H.2 CPP Evaluation Details & Results
@@
Appendix I Community Language Simulation Details
I.1 Community Language Simulation Prompts
RATING DEFINITIONS: ====================================================================================================== {{RATING DEFINITION}} ======================================================================================================
POST TITLE (context): {{POST TITLE}}
COMMENT: {{COMMENT BODY}}
REWRITTEN COMMENT:
RATING DEFINITIONS: =================================================================================================================================== "formality": """1. "Very Casual": extensive use of abbreviations, slangs, non-standard capitalization, missing syntactic components (no noun, no verb in sentence), incorrect punctuations, colloquialisms, contractions, inappropriate language (e.g. cuss words). 2. "Somewhat Casual": existence of slangs, missing syntactic components (no noun, no verb in sentence), unnecessary use of exclamation marks, inappropriate language (e.g. cuss words, “idiots”), or persistent presence of nonstandard capitalization, missing/incorrect punctuations, abbreviations, colloquialisms, contractions, nonstandard grammar and spelling. 3. "Neutral": Presence of a few nonstandard capitalization (e.g. not capitalized first letter of sentence), missing/incorrect punctuation, nonstandard grammar and spelling, abbreviation, colloquialisms, and relatively complete sentences. No slangs or emojis. 4. "Somewhat Formal": syntactically well structured, correct capitalization, complete sentences, correct punctuation, correct grammar. No abbreviations, no slang, no colloquialisms, can have acronyms and contractions. Ex. “I appreciate it. Thank you.” 5. "Very Formal": very structured thoughts and professional language, no abbreviations/slang/contractions/colloquialisms, grammatically correct. Contains structure in terms of the content (topic sentence, explanation, reasoning, etc). Ex. “I appreciate your guidance *insert details*”""", =================================================================OR================================================================ "supportiveness": """1. "Very Unsupportive": Aggressive, attacking the OP or others. Extremely rude, unreasonable, or even psycho. Outright judging that others are wrong/inferior. Using extremely inappropriate language. 2. "Somewhat Unsupportive": rude, unfriendly, disrespectful, promotes toxic behavior, leads to negative atmosphere. Will make a (normal) reader a little uncomfortable. Using inappropriate language. 3. "Neutral": neither supportive or toxic. Usually short texts like “Coffee and music” which doesn’t include any supportiveness or toxicity features 4. "Somewhat Supportive": respectful, constructive comments that have a positive outlook, not necessarily zealously supportive. Usually the commentator makes an effort to answer the question. 5. "Very Supportive": extremely positive, encouraging, promotes supportive & uplifting discussion. (e.g. omg i absolutely love this!!!!!)""", =================================================================OR================================================================ "sarcasm": """1. "Very Genuine": extremely sincere, honest, no implications. Profound or heartfelt messages. 2. "Somewhat Genuine": sincere and authentic, not lying. Includes subjective opinions that have enough content and context to judge as genuine (i.e. not a few words). E.g. some helpful advice. 3. "Neutral": Neither genuine nor sarcastic. Often includes short, objective answers (i.e. 1-3 words) that don’t imply anything. 4. "Somewhat Sarcastic": appears nice, but actual meaning is opposite to textual meaning and is often negative. Often an intention to be funny. 5. "Very Sarcastic": extreme ridicule or mockery, implicitly insulting. Exaggerated verbal irony.""", =================================================================OR================================================================ "politeness": """1. "Very Rude": disrespectful, demanding, offensive tone. E.g. “get the fuck out, shut up.” 2. "Somewhat Rude": not considering others feelings, imposing, generalizing without knowing the full context. E.g. judgy: “people like you would never…”, giving unsolicited advice: “Never …!” or comments that don’t really answer the question. Using exclamation/all caps when unnecessary. Often does not save their own or other’s face. 3. "Neutral": neither showing concern for others’ “face” nor being disrespectful. E.g. “you can do this…,”. Often includes comments that are straightforward but not rude. “bald-on record politeness” in politeness theory. 4. "Somewhat Polite": Making individuals feel good about themselves (appealing to positive face) or making the individuals feel like they haven’t been imposed upon/taken advantage of (appealing to negative face). in case of agreement: friendliness and camaraderie, compliments, common grounds; in case of disagreeing opinions: not assuming, not coercing, recognizing and addressing the hearer’s right to make his or her own decisions freely. (E.g. No offense but…, People usually…, I’m sure you know more than I do but…, replacing “I” and “you” with “people” or “we”). “positive politeness” and “negative politeness” in politeness theory. 5. "Very Polite": showing concern for others. give hints, give clues of association, presuppose, understate, overstate, use tautologies. Rely on the hearer to understand implications (e.g. I would do…, do you think you want to…) “Off-record politeness” in politeness theory.""", =================================================================OR================================================================ "humor": """1. "Very Serious": language and tone indicative of solemnity or earnestness, with a focus on conveying information or opinions with gravity and sincerity. Look for expressions of concern, absence of humor, and a straightforward communication style. 2. "Somewhat Serious": maintains a moderate level of seriousness, can include a mix of formal and informal language, occasional expressions of concern, and a balance between conveying important information or opinions with some degree of approachability. 3. "Neutral": not trying to be serious or humorous, or striking a balance between seriousness and humor. includes neutral expressions, and a versatile communication style adaptable to the context. 4. "Somewhat Humorous": incorporates humor or light-hearted language in a manner that enhances the discussion without detracting from its overall message. Can include humorous anecdotes, and playful expressions that contribute positively to the conversation. 5. "Very Humorous": primarily focuses on humor and entertainment, with language and expressions intended to amuse other users. Include witty remarks and humorous anecdotes that prioritize laughter and enjoyment over seriousness. ===================================================================================================================================
I.2 Filters for Community Language Simulation
@@
Appendix M Cost Estimation
M.1 GPT-3.5
M.2 GPT-4
[…]
Appendix N RPM Plots

### v2

@@ Appendix D Grounding 5-point Scale for Normness Ratings
In §4.2, we employ a 5-point Likert scale using GPT-3.5 to rate comments and sample them to gauge their normness. Additionally, in §4.2.2, we generate five variations of each original seed comment based on the 5 different scales of normness. Thus, for each norm dimension, we created a 5-point Likert scale and grounded their definitions in prior works Dementieva et al. (2022); Wulczyn et al. (2017); Joshi et al. (2017); Goffman (1955); Brown and Levinson (1987); Lakoff (1973). For example, we define formality based on using abbreviations, slang, colloquialisms, non-standard capitalizations, complete sentences, contractions, punctuations, and opening expressions of sentences Dementieva et al. (2022). Meanwhile, we define politeness as a set of strategies for conducting face-threatening acts while minimizing the chance that we or others will lose our positive or negative faces. Brown and Levinson (1987). The 5-point Likert scale across the norm dimensions can be found in Figures 61-63 as well as Figure 10.
Appendix E GPT Evaluations
Recall in §4.2.1 that we employ GPT-3.5 to sample and rate comments on a 5-point Likert scale (defined in Appendix D) for a particular norm dimension and subsequently use GPT-4 to generate binary synthetic labels comparing a pair of comments. In Appendix E.1, we describe the process of curating human annotations. In Appendix E.2, we evaluate the quality of GPT-3.5 rating. In Appendix E.3, we describe our prompt design considerations and prompt tuning results. In Appendix E.4, we evaluate the final GPT-4 automatic pairwise labeling pipeline using the human annotations.
@@ E.1 Normness Scale Annotation
To evaluate the NSP models and the quality of GPT-4 generated labels for student models, we curate a high-quality human annotation set of 450 samples for each norm dimension. The human annotations of norms are challenging due to subjectivity. To reduce subjectivity, we conducted training sessions with annotators and iteratively improved our annotation guidelines, grounding the definitions of various norms based on prior works (see Appendix D). Each sample was annotated by three volunteer annotators, who are graduate students in NLP and Linguistics at a US-based institution and familiar with the subreddits in our study. We did not provide payment, but we obtained consent to use their annotations for AI model evaluation.
@@
Across the four topics, we collected human annotations for 450 samples [^11] For all topics except “Gender,” we annotated 100 randomly-sampled pairwise comments. For “Gender” topic, we annotated 150 pairwise comments, in which 100 pairwise comments came from r/askmen and r/askwomen while the remaining 50 pairwise comments came from comparisons with one of the gender subreddits (including r/asktransgender) and r/asktransgender.. Each sample was annotated for five norm dimensions, resulting in a total of 2,250 annotations per human annotator.
The average inter-annotator agreement, measured by Fleiss’s $\kappa$, was 0.56, considered a moderate agreement Landis and Koch (1977). Due to the nuance and subtlety of norms, Fleiss’s $\kappa=0.56$ provides a solid foundation for our annotation labels. For instance, Passonneau and Carpenter (2014) reported scores as low as 0.2 in subjective tasks such as word sense annotations. Refer to Table 5 for the full agreement scores across 4 topics and 5 norm dimensions.
@@
E.2 Evaluating the Quality of GPT-3.5 Rating
@@
E.3 GPT-4 Automatic Pairwise Labeling
We underwent extensive prompt-tuning efforts to generate high-quality and accurate binary synthetic labels using GPT-4. Below, we discuss our prompt design choices (§E.3.1), the prompt tuning results to select the best prompt for our task (§E.3.2), and the full evaluation results of the chosen prompt against human annotations (§E.4).
E.3.1 Prompt Design Considerations
@@
E.3.2 Prompt-Tuning Results
Based on the proposed features in §E.3.1, we design multiple prompting pipelines and evaluate their performance on the binary labeling task—given two comments, compare the comments in each of the five norm dimensions. Performance is measured by the label accuracy against a human-annotated gold data, thus assessing the effect of different prompting pipelines to produce accurate labels.
@@ E.3.2 Prompt-Tuning Results
E.4 Evaluating the Chosen GPT-4 Labeling Pipeline
@@ G.3 NSP Evaluation Results
Appendix H Community Language Simulation Details
Here, we describe the details of the community language simulation (CLS). In Appendix H.1, we describe the CLS prompts to generate style-transferred comments that adopt the intended norm dimension (e.g., more sarcastic). In Appendix H.2, we describe our data filtering pipeline to ensure the quality of the synthetic comments. In Appendix H.3, we conduct human evaluation to validate the quality of the filtered synthetic comments across content preservation, fluency, naturalness, and overall quality. In Appendix H.4, we evaluate the faithfulness of the community language simulation; specifically, we validate whether the style-transferred comments adopted the intended norm dimension.
H.1 Community Language Simulation Prompts
RATING DEFINITIONS: ============================================================================= {{RATING DEFINITION}} =============================================================================
POST TITLE (context): {{POST TITLE}}
COMMENT: {{COMMENT BODY}}
REWRITTEN COMMENT:
RATING DEFINITIONS: ==================================================================================================
"formality": """1. "Very Casual": extensive use of abbreviations, slangs, non-standard capitalization, missing syntactic components (no noun, no verb in sentence), incorrect punctuations, colloquialisms, contractions, inappropriate language (e.g. cuss words). 2. "Somewhat Casual": existence of slangs, missing syntactic components (no noun, no verb in sentence), unnecessary use of exclamation marks, inappropriate language (e.g. cuss words, “idiots”), or persistent presence of nonstandard capitalization, missing/incorrect punctuations, abbreviations, colloquialisms, contractions, nonstandard grammar and spelling. 3. "Neutral": Presence of a few nonstandard capitalization (e.g. not capitalized first letter of sentence), missing/incorrect punctuation, nonstandard grammar and spelling, abbreviation, colloquialisms, and relatively complete sentences. No slangs or emojis. 4. "Somewhat Formal": syntactically well structured, correct capitalization, complete sentences, correct punctuation, correct grammar. No abbreviations, no slang, no colloquialisms, can have acronyms and contractions. Ex. “I appreciate it. Thank you.” 5. "Very Formal": very structured thoughts and professional language, no abbreviations/slang/contractions/colloquialisms, grammatically correct. Contains structure in terms of the content (topic sentence, explanation, reasoning, etc). Ex. “I appreciate your guidance *insert details*”""",
==================================================OR===============================================
"supportiveness": """1. "Very Unsupportive": Aggressive, attacking the OP or others. Extremely rude, unreasonable, or even psycho. Outright judging that others are wrong/inferior. Using extremely inappropriate language. 2. "Somewhat Unsupportive": rude, unfriendly, disrespectful, promotes toxic behavior, leads to negative atmosphere. Will make a (normal) reader a little uncomfortable. Using inappropriate language. 3. "Neutral": neither supportive or toxic. Usually short texts like “Coffee and music” which doesn’t include any supportiveness or toxicity features 4. "Somewhat Supportive": respectful, constructive comments that have a positive outlook, not necessarily zealously supportive. Usually the commentator makes an effort to answer the question. 5. "Very Supportive": extremely positive, encouraging, promotes supportive & uplifting discussion. (e.g. omg i absolutely love this!!!!!)""",
==================================================OR===============================================
"sarcasm": """1. "Very Genuine": extremely sincere, honest, no implications. Profound or heartfelt messages. 2. "Somewhat Genuine": sincere and authentic, not lying. Includes subjective opinions that have enough content and context to judge as genuine (i.e. not a few words). E.g. some helpful advice. 3. "Neutral": Neither genuine nor sarcastic. Often includes short, objective answers (i.e. 1-3 words) that don’t imply anything. 4. "Somewhat Sarcastic": appears nice, but actual meaning is opposite to textual meaning and is often negative. Often an intention to be funny. 5. "Very Sarcastic": extreme ridicule or mockery, implicitly insulting. Exaggerated verbal irony.""", ==================================================OR===============================================
"politeness": """1. "Very Rude": disrespectful, demanding, offensive tone. E.g. “get the fuck out, shut up.” 2. "Somewhat Rude": not considering others feelings, imposing, generalizing without knowing the full context. E.g. judgy: “people like you would never…”, giving unsolicited advice: “Never …!” or comments that don’t really answer the question. Using exclamation/all caps when unnecessary. Often does not save their own or other’s face. 3. "Neutral": neither showing concern for others’ “face” nor being disrespectful. E.g. “you can do this…,”. Often includes comments that are straightforward but not rude. “bald-on record politeness” in politeness theory. 4. "Somewhat Polite": Making individuals feel good about themselves (appealing to positive face) or making the individuals feel like they haven’t been imposed upon/taken advantage of (appealing to negative face). in case of agreement: friendliness and camaraderie, compliments, common grounds; in case of disagreeing opinions: not assuming, not coercing, recognizing and addressing the hearer’s right to make his or her own decisions freely. (E.g. No offense but…, People usually…, I’m sure you know more than I do but…, replacing “I” and “you” with “people” or “we”). “positive politeness” and “negative politeness” in politeness theory. 5. "Very Polite": showing concern for others. give hints, give clues of association, presuppose, understate, overstate, use tautologies. Rely on the hearer to understand implications (e.g. I would do…, do you think you want to…) “Off-record politeness” in politeness theory.""", ==================================================OR===============================================
"humor": """1. "Very Serious": language and tone indicative of solemnity or earnestness, with a focus on conveying information or opinions with gravity and sincerity. Look for expressions of concern, absence of humor, and a straightforward communication style. 2. "Somewhat Serious": maintains a moderate level of seriousness, can include a mix of formal and informal language, occasional expressions of concern, and a balance between conveying important information or opinions with some degree of approachability. 3. "Neutral": not trying to be serious or humorous, or striking a balance between seriousness and humor. includes neutral expressions, and a versatile communication style adaptable to the context. 4. "Somewhat Humorous": incorporates humor or light-hearted language in a manner that enhances the discussion without detracting from its overall message. Can include humorous anecdotes, and playful expressions that contribute positively to the conversation. 5. "Very Humorous": primarily focuses on humor and entertainment, with language and expressions intended to amuse other users. Include witty remarks and humorous anecdotes that prioritize laughter and enjoyment over seriousness. ==================================================================================================
H.2 Filters for Community Language Simulation
H.3 Community Language Simulation Filter Annotation
Topic Completely Share Roughly Mostly Completely
Dissimilar Details Equiv. Equiv. Equiv.
Gender 0.02 0.16 0.16 0.51 0.16
Politics 0.02 0.16 0.29 0.4 0.13
Science 0.04 0.07 0.13 0.42 0.33
Finance 0.03 0.1 0.12 0.38 0.37
Total 0.03 0.12 0.17 0.43 0.26
Topic Not at all Somewhat Very
Gender 0.04 0.16 0.80
Politics 0.04 0.09 0.87
Science 0.00 0.16 0.84
Finance 0.07 0.17 0.77
Total 0.04 0.14 0.82
Topic High Quality Not High Quality
Gender 0.67 0.33
Politics 0.58 0.42
Science 0.91 0.09
Finance 0.70 0.30
Total 0.71 0.29
Topic Original Comments Synthetic Comments
Gender 0.96 0.43
Politics 0.73 0.13
Science 0.75 0.78
Finance 0.42 0.78
Total 0.81 0.50
Model Formality Politeness Humor Supportiveness Sarcasm Verbosity Average
GPT-4o Judge 0.93 0.96 0.87 0.95 0.84 0.87 0.90
Normness model 0.88 0.84 0.76 0.75 0.64 0.91 0.80
Table 15: Evaluation results on the faithfulness of the community language simulation. We sampled 1,560 pairs of original and Llama3-8b-Instruct generated style-transferred comments (e.g. rewritten to be more sarcastic) and used GPT-4o as a judge to determine whether the comment is, for example, more sarcastic than the original one, finding an average percentage agreement of 90%. In addition, we checked whether the intended change by the prompt in the style transfer aligned with the normness scale predictor model, finding an average percentage agreement of 80% across topics and norm dimension.
H.4 Faithfulness of the Community Language Simulation
After conducting human evaluations to assess the content preservation, fluency, naturalness, and overall quality of the generated comments, we evaluated the faithfulness of the community language simulation. Specifically, we validated whether the style-transferred comments adopted the intended norm dimension (e.g., more sarcasm) when prompted to. To do this, we sampled 1,560 pairs of original and Llama3-8b-Instruct generated style-transfer comments and conducted two validations. Table 15 contains the validation results.
Our validations demonstrate that the style-transferred comments successfully adopted the intended norm dimension when prompted. First, we employed GPT-4o as a judge to determine whether the generated comment had, for instance, become more sarcastic than the original one, finding an average percentage agreement of 90%. Across the norm dimensions, we found that GPT-4o agreed with the intended style transfer, with percentage agreement rates ranging from 84%-96%. Second, we validated whether the intended change by the prompt in the style transfer aligned with the normness scale predictor model (NSP), finding an average percentage agreement of 80% across the topics and norm dimensions. These validations collectively indicate that the style-transferred comments effectively captured the intended shifts in the norm dimension (e.g., becoming sarcastic).
Appendix I Community Preference Prediction
I.1 CPP Training Details
@@
I.2 CPP Evaluation Details & Results
@@
Appendix M RPM Plots
@@
M.4 Science Subreddits
Figure 33: RPM plots for science subreddits on the sarcasm dimension.
Figure 34: RPM plots for science subreddits on the humor dimension.
Appendix N RPM Plots with Original Comments
RPM plots that feature only the original comments. Unlike standard RPM plots, which display the difference in normness scale and z-score between original and style-transferred comments, these plots use the absolute values: normness scale on the x-axis and z-score on the y-axis.
Although these RPM plots don’t show how changes in normness influence community approval, they do provide insight into the average normness of comments across various communities.
N.1 Gender Subreddits
Figure 37: RPM plots for gender subreddits on the formality dimension.
Figure 39: RPM plots for gender subreddits on the sarcasm dimension.
Figure 40: RPM plots for gender subreddits on the humor dimension.
Figure 41: RPM plots for gender subreddits on the verbosity dimension.
N.2 Finance Subreddits
Figure 45: RPM plots for finance subreddits on the sarcasm dimension.
Figure 46: RPM plots for finance subreddits on the humor dimension.
N.3 Politics Subreddits
Figure 51: RPM plots for politics subreddits on the sarcasm dimension.
Figure 52: RPM plots for politics subreddits on the humor dimension.